OpenStudy (anonymous):

can I get help please? -7(t-6)-(t+6)=9

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

distribute??

OpenStudy (anonymous):

multiply -7 to (t-6) and -1 to (t+6)

OpenStudy (anonymous):

-7(t)-(-7(6))-(t)-(6)=9

OpenStudy (anonymous):

Im not getting it!

OpenStudy (anonymous):

Expand the expression on the Left Hand Side\[-7(t-6)-(t+6)\] \[-7t+42-t-6=9\] collect the t's together and the numbers together \[-8t+36=9\] solve for t \[-8t=-27\] \[t=\frac{ -27 }{ -8 }\] which is then equal to \[t=3\frac{ 3 }{ 8 }\] So, all you have to do is to make the expression big (expand it), collect the letters and numbers together (collect like terms) and then solve for the letter (here we solved for "t") Hope that helped!

OpenStudy (anonymous):

Thank you. It really helped!